OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

Volume of cone \[\frac{ 1 }{ 3 }\pi ^{2}h\] What are you having trouble on?

OpenStudy (anonymous):

\[V=\frac{1}{ 3 }\pi r^{2}h\]

OpenStudy (anonymous):

i know the formula i was looking for an answer

OpenStudy (anonymous):

\[100=\frac{1}{3}\pi7^2*h\]

OpenStudy (anonymous):

h=?

OpenStudy (anonymous):

h=300/(49*pi)

OpenStudy (anonymous):

\[ 300=\frac{ 1 }{ 3 }(3.14)(9)^2h\]